. If in Problem 3 the plastic tube between bottle and needle has an inner diameter of 7 mm and a length of 1.5 m, what fractional contribution does this tube make to the resistance to flow in the transfusion? 


1
Expert's answer
2021-11-04T10:17:52-0400

The viscosity of blood "\eta=0.004\ Pa\cdot s"


"R_h=\Delta p/Q=8\eta\Delta x/(\pi r^4)="


"=8\cdot0.004\cdot1.5/(3.14\cdot0.0035^4)=101.8\cdot10^6\ (Pa\cdot s/m^2)"